# Exploring Remote Collaboration: The Impact of Avatar Representation on Dyadic Haptic Interactions in Shared Virtual Environments

### v1

@@ II-D Plate control task
The trajectory of the target followed a rose curve, which moved according to the progression of time. The movement of the target was calculated using Equation (4).
${\bm{x}^{tgt}}=r\left[\begin{matrix}\sin\left(\frac{4\pi\cdot t\operatorname{fmod}(t^{per})}{dt}\right)\cos\left(\frac{2\pi\cdot t\operatorname{fmod}(t^{per})}{dt}\right)\\ \sin\left(\frac{4\pi\cdot t\operatorname{fmod}(t^{per})}{dt}\right)\sin\left(\frac{2\pi\cdot t\operatorname{fmod}(t^{per})}{dt}\right)\end{matrix}\right].$ (4)
Here, ${\bm{x}^{tgt}}=\left[\begin{matrix}x^{tgt}&y^{tgt}\end{matrix}\right]^{T}$ represents the position vector of the target on the plate, indicating its x and y axes position. $t$is the time, and $t^{per}$ is the period of the rose curve from start to finish, designed to be 20 seconds in this study. $r$ is the radius of the trajectory, set to 0.2 m. Dyads were not shown the trajectory visually during the experiment but memorized it during practice sessions.
The ball moved according to the plate’s inclination, governed by the dynamics applied in Equation (5).
@@
Here, ${\bm{x}^{bll}}=\left[\begin{matrix}x^{bll}&y^{bll}\end{matrix}\right]^{T}$ is the position vector of the ball on the plate. $\dot{\bm{x}}^{bll}$ is the velocity vector of the ball. $m$ is the ball mass, $g$ is the gravitational acceleration, $\eta$ is the viscosity coefficient, with values set to 0.014, 9.8, and 1.5, respectively, in this study. $m$ and $\eta$ are critical parameters determining the movement of the ball relative to the tilt of the plate, significantly influencing the difficulty of the task. These parameters were chosen experimentally to set the difficulty of the task.
II-E Co-localization
@@
The landmark was also used to share the positions of avatars. When a participant grasped the plate, it was essential that the avatar of their partner appeared to be grasping the plate as well, enhancing the sense of collaboration and presence within the SVE. The avatar’s position was captured from the landmark’s local coordinate system and transmitted to other devices, ensuring that the dyad could share the space around the plate completely, irrespective of the interface’s physical location. The avatar position data were updated and transmitted every 10 milliseconds, allowing participants to focus on their tasks without experiencing any perceivable delay. This technology facilitated a seamless integration of physical and virtual interactions, crucial for maintaining high levels of engagement and performance in collaborative virtual tasks.
III EXPERIMENT
This paper investigates the impact of avatar representation in remote physical collaboration. We hypothesized that haptic interaction alone could sustain social presence among dyads, even without the visual representation of avatars. To test this, we conducted an experiment in which dyads performed a plate control task through haptic interaction while varying the avatar representation of the participants and their partners. Previous research indicates that the absence of an avatar significantly reduces social presence [8]. Consequently, the variation in avatar representation was limited to visible or invisible in our experiment.
III-A Experimental design
  • ALL: Both participant’s and partner’s avatars were displayed in the virtual environment. It assumed that both users had body tracking devices.
  • OTHER: Only the partner’s avatar was displayed.
The conditions ALL, ALONE, and NONE are those anticipated in real environments. On the other hand, OTHER has been newly added based on our previous research [38], but it is a condition that does not occur in reality (as only the partner receives body tracking information, and no avatar is displayed on one’s own screen). However, we were interested in the effects that occur when self-avatar is not displayed. If self-avatar alone is not displayed in the SVE, can participants maintain their sense of social presence? In addition, each condition was applied simultaneously to both participants.
III-B Research hypothesis
Haptic interaction and visual interaction through avatar representation contribute to providing the presence of a partner. Even without avatar representation, participants may still feel the presence of the other through haptic interaction, and the sense of social presence may not be lost. On the other hand, we speculate that dyads will work hard to achieve the objectives of the task, regardless of the conditions. Therefore, we propose the following hypotheses:
  • H1: Haptic interaction continues to provide the presence of a partner, even when avatars are not displayed. Thus, the social presence remains the same across all conditions.
  • H2: Regardless of avatar representation, the dyads will continue to achieve the best performance possible. Thus, the performance is consistent across all conditions.
  • H3: Changes in avatar representation can psychologically affect participants and alter the intensity of their efforts. Thus, there is a variation in the intensity of the effort between conditions.
III-C Participants
28 participants were included in this experiment, all university students. Participants provided informed consent according to the Declaration of Helsinki before the experiment. The experiment was approved by the Human Life Ethics Committee of Tokyo Denki University. Among the participants, 24 were right-handed and four left-handed. The group included 24 males and four females aged 21 to 28 ($M=22.96$, $SD=2.08$).
@@ III-D Procedure
Initially, participants provided personal information (biological sex, age, height, and dominant hand) to collect physical data. The participants then practiced alone to familiarize themselves with the plate control task, SVE, and HMD. They wore the Meta Quest 3 and calibrated it to synchronize the haptic interface with the virtual environment, verifying that their avatar tracked their body, hands, and fingers accurately. During practice, participants could ask the experimenter questions about the plate control task without restrictions. The practice continued until participants declared comfortable with the plate control task, with no time limits set by the experimenter.
@@
After practice, each dyad collaboratively performed the plate control task for 80 seconds. The avatar representation condition was applied randomly from four conditions, and visual interaction between the dyads was presented. During the experiment, verbal communication between the dyads was prohibited by the experimenter. Following the plate control task, participants removed the Meta Quest 3 and completed a questionnaire as instructed by the experimenter (details in III-E1). The questionnaire was displayed on a computer. After completing the questionnaire, the participants were given at least a two-hour break to consider fatigue and memory of their responses. After the break, participants performed the plate control task under a different condition and completed the questionnaire again. The experimenter conducted the plate control task and questionnaire four times with each pair to compare the four conditions.
III-E Measures
@@
A systematic review on the effects of avatars and agents during the use of HMD highlights that avatars can be effective or ineffective, depending on the task and situation [19]. They also point out that combining subjective evaluations with numerical data enhances the validity of research findings. Bailenson et al. argue that subjective measures, self-reported, depend on the participant’s ability to accurately assess their experiences [41]. Subjective measures are susceptible to participant bias, which can affect reliability. However, objective measures can detect responses to digital representations, such as personal distance, which are not captured by self-reports, suggesting higher sensitivity. Thus, in this paper, user experience (social presence) is subjectively measured, while task-related metrics (performance and effort) are objectively measured.
III-E1 Subjective measures
@@
Social presence is the psychological sense that another person is “being there” when using media or communication technologies [42]. This concept delivers the intimacy and immediacy of real interactions despite physical separation. A higher richness of information in communication media can create a stronger sense of social presence. To deepen understanding of how people feel connected in virtual environments, social presence is often used [7].
@@
To measure social presence, this study used the Networked Minds Social Presence Questionnaire (NMSPQ) [43, 44]. The NMSPQ includes five sub-scales: Co-presence, Perceived Attentional Engagement, Perceived Emotional Contagion, Perceived Comprehension, and Perceived Behavioral Interdependence. Since the interaction in our task was limited to non-verbal communication, participants had no opportunity for conversation and could not comprehend their partner’s thoughts. In addition, the general avatars used were expressionless and lacked personality, which prevented the transmission of emotions to the partner. For these reasons, Perceived Comprehension and Perceived Emotional Contagion were removed from the questionnaire.
@@
III-E2 Objective measures
In this paper, performance and effort were used as objective measures. The objective of the plate control task was to move a ball on a target, which required the dyads to collaborate to decrease the distance between the target and the ball. We decided to measure performance by calculating the root-mean-squared error (RMSE) between the target and the ball. Performance is defined as follows:
@@ III-E2 Objective measures
Effort was also measured. Although participants applied force to move the plate, not all applied force effectively contributed to the movement. According to Newton’s laws, when two equal forces are applied in opposite directions to an object, they cancel out, and the object does not move. In contrast, if forces are applied in the same direction, the object moves. Thus, the difference between the total forces applied to the object and the forces that actually moved the object represents the effort not contributing to the movement, considered as effort here. Although effort does not directly impact performance, it is considered crucial for haptic communication used by dyads to negotiate for achieving high performance.
The total force applied to the plate is given by:
@@
Here, $k$ represents the number of steps ($k=1\ldots N$), $L$ is the half length of the plate used to convert torque into force. This measure helps to understand the extent of unproductive effort exerted by participants during the task.
@@ IV-A Subjective measures
[…]
@@
SP CP PAE PBI
MD pbonf MD pbonf MD pbonf MD pbonf
ALL ALONE $1.148$ $<$ .001 $1.612$ $<$ .001 $1.179$ $<$ .001 $0.500$ $0.147$
OTHER $0.489$ $0.062$ $0.728$ $0.035$ $0.304$ $1.000$ $0.357$ $0.633$
NONE $1.448$ $<$ .001 $2.143$ $<$ .001 $1.071$ $<$ .001 $0.899$ $<$ .001
ALONE OTHER $-0.659$ $0.004$ $-0.884$ $0.005$ $-0.875$ $0.001$ $-0.143$ $1.000$
NONE $0.300$ $0.669$ $0.531$ $0.250$ $-0.107$ $1.000$ $0.399$ $0.427$
OTHER NONE $0.959$ $<$ .001 $1.415$ $<$ .001 $0.768$ $0.005$ $0.542$ $0.091$
Note. P-value adjusted for comparing a family of 6. MD stands for Mean Difference.
SP, CP, PAE and PBI stand for Social Presence, Co-presence, Perceived Attentional Engagement
and Perceived Behavioral Interdependence, respectively.
Figure 3: Subjective measures of Social Presence (NMSPQ). From left to right, Social Presence, Co-presence, Perceived Attentional Engagement and Perceived Behavioral Interdependence.
@@ IV-B Objective measures
[…]
Given this nonnormality, the Friedman test, a nonparametric equivalent to repeated measures ANOVA, was performed to check for differences across the conditions ALL, ALONE, OTHER, and NONE in terms of Performance and Effort. The results showed that there were no significant differences in either performance or effort under the conditions. Comparisons of performance and effort under the different conditions are illustrated in Fig. 4 and Fig. 5, respectively.
@@
Additionally, changes in performance were observed over time. Fig. 6 shows a comparison of the temporal changes in performance. Here, a ’Quarter Period’ represents the average of each quarter of a rose curve, with the period of the rose curve being 20 s (i.e. the one-quarter period is 5 s). The plate control task was conducted for 80 seconds, so the graph displays 16 quarter periods. Performance for each quarter period was normalized based on performance in the first quarter period.
As a result, it was confirmed that the dyads improved their performance compared to the first quarter period under ALL, ALONE, OTHER and NONE conditions (lower values indicate better performance). However, no clear differences were detected between the conditions ALL, ALONE, OTHER and NONE. This suggests that while performance improved during the cooperative task, the specific condition did not significantly influence the rate or extent of improvement.
@@ V-A Subjective measures
The survey revealed significant differences in all measures of Social Presence, Co-presence, Perceived Attentional Engagement, and Perceived Behavioral Interdependence. Notably, significant differences were found in comparisons of the co-presence sub-scale, except between the ALONE and NONE conditions. These results led to the rejection of the hypothesis ${\bf H1}$.
@@
Social presence scores in the OTHER condition were the next highest after ALL, but ALONE and NONE showed lower scores compared to ALL and OTHER. This suggests that having a self-avatar is less critical for maintaining a high social presence than seeing the partner’s avatar. Participants were constantly engaged in collaborative tasks that involved plate handling during the experiments, a constant factor in all avatar conditions. It is clear that the partner’s presence was sensed through haptic interaction regardless of the avatar’s visibility. However, the absence of the partner’s avatar reduced the social presence. Haptic feedback plays a vital role in SVEs, and several studies have reported that it improves social presence [14, 15, 16, 17, 18]. These results imply that visual representation of a partner, such as avatars, might be more critical than haptic interaction for enhancing social presence.
Next, we focus on the perceived behavioral Interdependence. Our results indicate that the ALL condition scored significantly higher than NONE. Perceived Behavioral Interdependence measures the extent to which users’ actions affect and are affected by their partners. Thus, participants felt their influence on or influence from the partner was more significant when avatars were visible. This finding is supported by studies such as those by Heidicker [9] and Yoon [45]. However, in our study, the dyads were constantly provided with haptic feedback, and their influence on each other always existed between the dyads, regardless of avatar visibility. Yet, despite this, avatars influenced participants, potentially feeling monitored by their partners (similar to how one might behave more cautiously when a camera is on during a web conference). This observation suggests that it may be possible to choose to omit the display of avatars in order to reduce the mental burden on users.

### v2

@@ III-C4 Plate control task
The trajectory of the target followed a rose curve, which moved according to the progression of time. The movement of the target was calculated as:
${\bm{x}^{tgt}}=r\left[\begin{matrix}\sin\left(4\pi\zeta\right)\cos\left(2\pi\zeta\right)\\ \sin\left(4\pi\zeta\right)\sin\left(2\pi\zeta\right)\end{matrix}\right].$ (4)
Here, ${\bm{x}^{tgt}}=\left[\begin{matrix}x^{tgt}&y^{tgt}\end{matrix}\right]^{T}$ represents the position vector of the target on the plate, indicating its x and y axes position. $\zeta=\operatorname{fmod}\left(t,t^{per}\right)/t^{per}$ is the quotient remainder of $t$ and $t^{per}$ divided by $t^{per}$, which is a ratio of 0.0 to 1.0. This can be computed by using the fmod function in C, Pyhon, etc. $t$ represents time, and $t^{per}$ is the period of the rose curve from start to finish, designed to be 20 s in this study. $r$ is the radius of the trajectory, set to 0.2 m. Dyads were not shown the trajectory visually during the experiment but memorized it during practice sessions.
The ball moved according to the plate’s inclination, governed by the dynamics applied as:
@@
Here, ${\bm{x}^{bll}}=\left[\begin{matrix}x^{bll}&y^{bll}\end{matrix}\right]^{T}$ is the position vector of the ball on the plate. $\dot{\bm{x}}^{bll}$ is the velocity vector of the ball. $m$ is the ball mass, $g$ is the gravitational acceleration, $\eta$ is the viscosity coefficient, with values set to 0.014, 9.8, and 1.5, respectively, in this study. $m$ and $\eta$ are critical parameters determining the movement of the ball relative to the tilt of the plate, thereby significantly influencing the difficulty of the task. These parameters were chosen experimentally to set the difficulty of the task.
III-C5 Co-localization
@@
The landmark was used to share the positions of avatars. When a participant grasped the plate, it was essential that the avatar of their partner appeared to be grasping the plate as well, thereby enhancing the sense of collaboration and presence within the SVE. The avatar’s position was captured from the landmark’s local coordinate system and transmitted to other devices, ensuring that the dyad could share the space around the plate completely, irrespective of the interface’s physical location. The avatar position data were updated and transmitted every 10 milliseconds, allowing participants to focus on their tasks without experiencing any perceivable delay. This technology facilitated a seamless integration of physical and virtual interactions, crucial for maintaining high levels of engagement and performance in collaborative virtual tasks.
III-D Participants
We performed a priori power analysis using G*Power [44] to determine our sample size. To detect a medium effect size, we selected an effect size of f = 0.25, an alpha error probability of 0.05, and a power of 0.8. Because we planned to conduct a repeated measures ANOVA with four experimental conditions, we set the number of groups to one while of measurements to four. Consequently, 28 participants were included in this experiment, all university students. Participants provided informed consent according to the Declaration of Helsinki before the experiment. The experiment was approved by the Human Life Ethics Committee of Tokyo Denki University. Among the participants, 24 were right-handed and four left-handed. The group included 24 males and 4 females aged 21 to 28 ($M=22.96$, $SD=2.08$).
@@ III-E Procedure
Initially, participants provided personal information (biological sex, age, height, and dominant hand) to collect physical data. The participants practiced alone to familiarize themselves with the plate control task, SVE, and HMD. They wore the Meta Quest 3 and calibrated it to synchronize the haptic interface with the VE, verifying that their avatar tracked their body, hands, and fingers accurately. During practice, participants could ask the experimenter questions about the plate control task without restrictions. The practice continued until the participants declared that they were comfortable with the plate control task, with no time limits set by the experimenter.
@@
After practice, each dyad collaboratively performed the plate control task for 80 s. The avatar representation condition was applied randomly from four conditions, and visual interaction between the dyads was presented. During the experiment, verbal communication between the dyads was prohibited by the experimenter. Following the plate control task, participants removed the Meta Quest 3 and completed a questionnaire as instructed by the experimenter (details in III-F1). The questionnaire was displayed on a computer. Because we employed a within-subjects design, all participants were required to experience each of the four conditions.
After completing the questionnaire, the participants were provided with at least a two-hour rest to account for fatigue and memory of their responses. On each day of participation, they completed two of the four experimental conditions, thus, they requiring two days to finish all conditions. Consequently, each participant spent approximately four hours per day in the experiment, that is, eight hours over the two days. No restrictions were placed on participants’ activities during the test. After the rest, participants performed the plate control task under a different condition and completed the questionnaire again. The experimenter conducted the plate control task and questionnaire four times with each pair to compare the four conditions.
III-F Measures
@@
A systematic review was conducted on the effects of avatars and agents during the use of HMD highlights that avatars can be effective or ineffective, depending on the task and situation [19]. Furthermore, they also point out that combining subjective evaluations with numerical data enhances the validity of research findings. Bailenson et al. argue that subjective measures, self-reported, depend on the participant’s ability to accurately assess their experiences [47]. Subjective measures are susceptible to participant bias, which can affect reliability. However, objective measures can detect responses to digital representations, such as personal distance, which are not captured by self-reports, suggesting higher sensitivity. Thus, in this study, user experience (social presence) is subjectively measured, while task-related metrics (performance and effort) are objectively measured.
III-F1 Subjective measures
@@
Social presence is the psychological sense that another person is “being there” when using media or communication technologies [48]. This concept delivers the intimacy and immediacy of real interactions despite physical separation. A higher richness of information in communication media can create a stronger sense of social presence. To deepen understanding of how people feel connected in VEs, social presence is often used [7].
@@
To measure social presence, this study used the Networked Minds Social Presence Questionnaire (NMSPQ) [49, 50]. The NMSPQ includes five sub-scales: Co-presence, Perceived Attentional Engagement, Perceived Emotional Contagion, Perceived Comprehension, and Perceived Behavioral Interdependence. Because the interaction in our task was limited to non-verbal communication, participants had no opportunity for conversation and could not comprehend their partner’s thoughts. In addition, the general avatars used were expressionless and lacked personality, which prevented the transmission of emotions to the partner. Therefore, Perceived Comprehension and Perceived Emotional Contagion were removed from the questionnaire.
@@
III-F2 Objective measures
In this study, performance and effort were used as objective measures. The objective of the plate control task was to move a ball on a target, which required the dyads to collaborate to decrease the distance between the target and the ball. We decided to measure performance by calculating the root-mean-squared error (RMSE) between the target and the ball. Performance is defined as follows:
@@ III-F2 Objective measures
Effort was measured. Although participants applied force to move the plate, not all applied force effectively contributed to the movement. According to Newton’s laws, when two equal forces are applied in opposite directions to an object, they cancel out, and the object does not move. However, if forces are applied in the same direction, the object moves. Thus, the difference between the total forces applied to the object and the forces that actually moved the object represents the effort not contributing to the movement, considered as effort here. Although effort does not directly impact performance, it is considered crucial for haptic communication used by dyads to negotiate for achieving high performance.
The total force applied to the plate is denoted by:
@@
Here, $k$ represents the number of steps ($k=1\ldots N$), and $L$ is the half length of the plate used to convert torque into force. This measure helps to understand the extent of unproductive effort exerted by participants during the task.
@@ IV RESULTS
Although participants performed the tasks in dyads, we treated each individual’s data as an independent sample for our analyses. This study aimed to compare individual evaluations and responses, rather than to analyze interactions within each dyad. Therefore, we did not employ a nested (multi-level) analysis that accounts for dyadic structure and instead performed a standard ANOVA to examine main effects. However, owing to the nature of the objective measures, we could not separate the data by individual; hence, we used one dataset per dyad for those measures.
Because participants could potentially observe each other’s faces and be influenced by sound or smell, there was a possibility of mutual interference. Nevertheless, our tasks and metrics were designed to be evaluated on an individual basis, and we did not anticipate systematic interactions sufficiently strong to bias the results (e.g., one participant’s responses heavily shaping the other’s). A preliminary examination of the data revealed no notable correlations or clustering among participants, reinforcing our decision not to model the dyad factor.
In the box plots presented in this study, the central line indicates the median, and the box edges represent the first and third quartiles. The whiskers show the range excluding outliers, determined using 1.5 times the interquartile range (IQR).
@@ IV-A Subjective measures
Subsequently, repeated measures ANOVA was conducted to compare scores under the four conditions (ALL, ALONE, OTHER, and NONE) for each scale. Mauchly’s test of sphericity indicated violations for Social Presence (p = .009) and Co-presence (p = .002), but not for Perceived Attentional Engagement (p = .994) or Perceived Behavioral Interdependence (p = .085). Consequently, the Greenhouse–Geisser correction was applied for Social Presence and Co-presence. The repeated measures ANOVA revealed significant main effect in Social Presence (F(2.296, 61.985) = 24.442, p $<$ .001, $\eta^{2}_{p}$ = .475), Co-presence (F(2.069, 55.858) = 27.301, p $<$ .001, $\eta^{2}_{p}$ = .503), Perceived Attentional Engagement (F(3, 81) = 13.404, p $<$ .001, $\eta^{2}_{p}$ = .332), and Perceived Behavioral Interdependence (F(3, 81) = 5.807, p = .001, $\eta^{2}_{p}$ = .177). To compare all group pairs, post hoc tests with Bonferroni-corrected p-values were performed. Detailed post hoc results are presented in Table I and Fig. 3.
@@
Pair MD SE t d pbonf
Social Presence
ALL ALONE 1.148 .186 6.158 1.521 $<$ .001
OTHER 0.489 .186 2.624 0.648 .062
NONE 1.448 .186 7.764 1.918 $<$ .001
ALONE OTHER -0.659 .186 -3.534 -0.873 .004
NONE 0.300 .186 1.609 0.397 .669
OTHER NONE 0.959 .186 5.143 1.270 $<$ .001
Co-presence
ALL ALONE 1.612 .257 6.281 1.604 $<$ .001
OTHER 0.728 .257 2.836 0.724 .035
NONE 2.143 .257 8.351 2.133 $<$ .001
ALONE OTHER -0.884 .257 -3.445 -0.880 .005
NONE 0.531 .257 2.070 0.529 .250
OTHER NONE 1.415 .257 5.515 1.409 $<$ .001
Perceived Attentional Engagement
ALL ALONE 1.179 .223 5.287 1.222 $<$ .001
OTHER 0.304 .223 1.362 0.315 1.000
NONE 1.071 .223 4.807 1.111 $<$ .001
ALONE OTHER -0.875 .223 -3.926 -0.907 .001
NONE -0.107 .223 -0.481 -0.111 1.000
OTHER NONE 0.768 .223 3.445 0.796 $<$ .005
Perceived Behavioral Interdependence
ALL ALONE 0.500 .218 2.292 0.534 .147
OTHER 0.357 .218 1.637 0.381 .633
NONE 0.899 .218 4.120 0.959 $<$ .001
ALONE OTHER -0.143 .218 -0.655 -0.152 1.000
NONE 0.399 .218 1.828 0.426 0.427
OTHER NONE 0.542 .218 2.483 0.578 0.091
Note. MD, SE, t, d and pbonf stand for Mean Difference, Standard Error, t-value, Cohen’s d and p-value with Bonferroni correction, respectively. p-value adjusted for comparing a family of 6.
Figure 3: Subjective measures of Social Presence (NMSPQ). From left to right, Social Presence, Co-presence, Perceived Attentional Engagement and Perceived Behavioral Interdependence (*: p $<$ 0.05, **: p $<$ 0.01, ***: p $<$ 0.001).
@@ IV-B Objective measures
For performance and effort, the Shapiro–Wilk test was performed for each condition (ALL, ALONE, OTHER, NONE). Performance showed a significant deviation from normality (p = .009), whereas effort did not (p = .796). Accordingly, the Friedman test (a nonparametric equivalent to repeated measures ANOVA) was used to assess whether performance differed across the four conditions. No significant main effect were identified ($\chi^{2}$(3) = 7.629, p = .054, W = .182).
Because effort met normality assumptions, repeated measures ANOVA was employed to test for differences across conditions. Mauchly’s test of sphericity for effort was no significant (p = .514). The ANOVA indicated no significant main effect (F(3, 39) = 1.375, p = .265, $\eta^{2}_{p}$ = .096). Comparisons of performance and effort under the different conditions are illustrated in Fig. 4 and Fig. 5, respectively.
@@
Additionally, changes in performance were observed over time. Fig. 6 shows a comparison of the temporal changes in performance. Here, a “Quarter Period” represents the average of each quarter of a rose curve, with the period of the rose curve being 20 s (i.e. the one-quarter period is 5 s). The plate control task was conducted for 80 s, thus the graph displays 16 quarter periods. Performance for each quarter period was normalized based on performance in the first quarter period.
Consequently, it was confirmed that the dyads improved their performance compared to the first quarter period under ALL, ALONE, OTHER, and NONE conditions (lower values indicate better performance). However, no clear differences were detected between the conditions ALL, ALONE, OTHER, and NONE. Therefore, while performance improved during the collaborative task, the specific condition did not significantly influence the rate or extent of improvement.
@@ V-A Subjective measures
The survey revealed significant main effect in all measures of Social Presence, Co-presence, Perceived Attentional Engagement, and Perceived Behavioral Interdependence. Notably, significant differences were found in comparisons of the co-presence sub-scale, except between the ALONE and NONE conditions. These results caused the rejection of the hypothesis ${\bf H1}$.
@@
Social presence scores in the OTHER condition were the next highest after ALL, but ALONE and NONE showed lower scores compared to ALL and OTHER. Therefore, having a self-avatar is less critical for maintaining a high social presence than seeing the partner’s avatar. Participants were constantly engaged in collaborative tasks that involved plate handling during the experiments, a constant factor in all avatar conditions. The partner’s presence was sensed through haptic interaction regardless of the avatar’s visibility. However, the absence of the partner’s avatar reduced the social presence. Haptic feedback plays a vital role in SVEs, and several studies have reported that it improves social presence [14, 15, 16, 17, 18]. These results indicate that visual representation of a partner, such as avatars, might be more critical than haptic interaction for enhancing social presence.
Subsequently, we focus on the perceived behavioral Interdependence. Our results indicate that the ALL condition scored significantly higher than NONE. Perceived Behavioral Interdependence measures the extent to which users’ actions affect and are affected by their partners. Thus, participants felt their influence on or influence from the partner was more significant when avatars were visible. This finding is supported by studies, such as those by Heidicker et al. [9] and Yoon et al. [51]. In our study, the dyads were constantly provided with haptic feedback, and their influence on each other always existed between the dyads, regardless of avatar visibility. However, avatars influenced participants, potentially feeling monitored by their partners (similar to how one might behave more cautiously when a camera is on during a web conference). This observation suggests that it may be possible to choose to omit the display of avatars to reduce the mental burden on users.